# Solitons in binary compounds with stacked two-dimensional honeycomb lattices

### v1

@@ VIII The role of interlayer disorder
For ${}^{\prime}$ stacking and a smooth soliton in the absence of disorder, Fig 12(x), the density of states shows a narrow peak at zero energy due to the flat bands related to the soliton states. Figure 16 shows the disorder-averaged density of states $\langle g(E)\rangle$ for a smooth soliton at the center $(8,8)$ of a system with ${}^{\prime}$ stacking and $N=16$ layers. Sharp disorder, Fig. 16(a) and (b), tends to destroy the narrow peak at zero energy, leaving a small but finite density of states in the bulk band gap. The narrow peak is fairly robust to smooth disorder, however, and more robust to hopping disorder than onsite disorder, Fig. 16(c) and (d). This is possibly because the solitons are textures in the onsite energies, not the hopping parameters.
IX Relevance to specific materials
@@
Our results apply to binary compounds with a honeycomb structure. A well-studied example is hexagonal boron nitride (h-BN) [27] which occurs predominantly with ${}^{\prime}$ stacking [57, 58, 60, 59, 61, 30, 31, 29, 62, 63], although samples with ABA [64, 31, 65, 66, 29, 67, 68] and ABC stacking [69, 70, 71, 72, 73] have also been fabricated. Density functional theory [28, 33, 30, 29] predicts that ABA stacking has a total energy comparable to ${}^{\prime}$, but that ABC stacking is less energetically favorable. Excitons localized on structural defects in h-BN have been discussed in the context of observed optical properties [74, 75, 76, 77], and features in X-ray spectroscopy have been attributed to interlayer stacking faults [78]. Domain walls in the in-plane direction have been observed experimentally [31, 79, 80] and discussed theoretically [81], and the nonlinear optical properties of multilayers with a controllable twist in the vertical (stacking) direction have been measured and analyzed [82].
The stacking types to be realized experimentally (${}^{\prime}$, ABA and ABC) all have different atoms connected in the vertical direction, i.e., boron to nitrogen, and a fault consisting of a connection between two identical atoms, i.e., boron to boron or nitrogen to nitrogen, will be energetically expensive. Single layer defects, as described in Section VII, were modeled using DFT [33] for AA, ${}^{\prime}$, ABA and ${}^{\prime}$ stacking, and it was found that the formation energy of such faults is very small. For example, an AA-type defect in ${}^{\prime}$ stacking was estimated to cost of the order of 505050\,50meV. Even so, it is likely that such a defect is difficult to realize experimentally because a relative shift in the in-plane direction would yield the more favorable ABA stacking [64, 29]. Nevertheless, experiments [83, 84] have recently demonstrated the possiblilty of engineering stacking of boron nitride (to create ferroelectric materials), and a relative twist between layers creates domains of ABA stacking separated, in the in-plane direction, by domain walls and regions of AA stacking.
@@
X Conclusions
Using a minimal tight-binding model, we determined the electronic properties of thin films of binary compounds with two atoms per unit cell arranged as stacked two-dimensional honeycomb lattices. The two atoms per cell are assigned different onsite energies. We considered six different stacking orders to determine whether a fault in the texture of onsite energies in the stacking direction supports localized states. Faults for ABC, ABA, and ${}^{\prime}$ stacking support localized states within the band gap, whereas faults for ${}^{\prime}$, ${}^{\prime}$, and AA stacking do not. For ABC and ABA stacking, there is a single localized state within the bulk band gap and, for ABC stacking, this state may hybridize with surface states. For ${}^{\prime}$ stacking, there are two states within the bulk band gap which may hybridize with each other, depending on parameter values.
We consider smooth solitons where the texture of onsite energies changes over a length scale greater than the interlayer spacing, leading to different band structures as compared to atomically sharp solitons. In particular, a smooth soliton in ${}^{\prime}$ stacking results in a flat band at zero energy and a corresponding narrow peak in the density of states. We show that this feature is fairly robust to long-range correlated disorder. Finally, we also consider the band structure due to single layer defects [33] where the signs of the onsite energies are reversed on a single layer only without a subsequent change of texture. Overall, different stackings and types of fault produce a range of contrasting electronic band structures and densities of states which will manifest themselves in various transport and spectroscopic measurements [74, 75, 76, 77, 78, 31, 79, 80].

### v2

@@ VIII The role of interlayer disorder
For AA stacking and a smooth soliton in the absence of disorder, Fig 12(x), the density of states shows a narrow peak at zero energy due to the flat bands related to the soliton states. Figure 16 shows the disorder-averaged density of states $\langle g(E)\rangle$ for a smooth soliton at the center $(8,8)$ of a system with AA stacking and $N=16$ layers. Sharp disorder, Fig. 16(a) and (b), tends to destroy the narrow peak at zero energy, leaving a small but finite density of states in the bulk band gap. The narrow peak is fairly robust to smooth disorder, however, and more robust to hopping disorder than onsite disorder, Fig. 16(c) and (d). This is possibly because the solitons are textures in the onsite energies, not the hopping parameters.
IX Electronic transport
Figure 17: Rhombohedral stacking with $N=10$ layers and onsite energy $U=0.1\,$eV showing band structures (left) and electrical conductivity (right). (a), (b) show the faultless system and (c), (d) show a system with an atomically-sharp soliton at the center using the minimal model with $\gamma_{2}=\gamma_{3}=\gamma_{4}=0$. (e), (f) show the faultless system and (g), (h) show a system with an atomically-sharp soliton at the center using the full parameter model. Parameter values are $\gamma_{0}=3.16\,$eV, $\gamma_{1}=0.381\,$eV, $\gamma_{2}=-0.02\,$eV, $\gamma_{3}=0.315\,$eV, $\gamma_{4}=0.044\,$eV, $a=2.46$Å. The conductivity calculation is performed for a system of length $L=1754a$ and width $W=1754a$, where $a$ is the lattice constant. Details are described in Appendix A.
Solitons, both atomically-sharp and smooth in position space, are able to support localized states with energies within the bulk band gap for ABC, ABA and AA stacking. The contrasting electronic band structures and densities of states will impact transport and spectroscopic measurements [51, 52, 53, 54, 55, 31, 56, 57]. As an example, we model coherent, ballistic, electronic transport for rhombohedral-stacking, generalizing a model developed for monolayer and bilayer graphene [58, 42] in which the central sample is connected to leads of the same material that are heavily doped in order to provide a large density of states. The conductivity is determined using the Landauer-Büttiker formalism [59] where the transmission probability is found by wave-matching at the boundaries of the leads to the central sample, and Hamiltonians are taken in the continuum limit. Details of our numerical calculations are described in Appendix A.
Generalizing previous results for monolayer and bilayer graphene [58, 42], we find that the minimal conductivity for rhombohedral stacking in the absence of onsite energies ($U=0$) is given by $\sigma_{\mathrm{min}}=4Ne^{2}/(\pi h)$ for $N$ layers with $W\gtrsim L$ where $W$ and $L$ are the sample width and length, respectively. Figure 17 shows band structures and corresponding conductivities for a rhombohedrally-stacked system with $N=10$ layers and onsite energy $U=0.1\,$eV. The band structures are plotted near the K point, with the direction of the $\Gamma$ and M points indicated for negative and positive $k_{x}$, respectively. Fig. 17(a) and (b) are for the minimal tight-binding model (with $\gamma_{0}$, $\gamma_{1}$, and $U$ only, Eq. (1)) and the faultless system. The region of the band gap, Fig. 17(a), corresponds to zero conductivity, Fig. 17(b), and, when the energy is equal to that of higher-energy bands, the presence of additional transport modes produces steps in the conductivity. Fig. 17(c) and (d) are for the minimal tight-binding model with an atomically-sharp soliton at the center. The state localized on the soliton creates a non-zero density of states resulting in non-zero conductivity, and the position of the strong hybridization of the soliton state with the surface states corresponds to a characteristic dip in the conductivity (at $E\approx U=0.1\,$eV in Fig. 17(d)).
X Additional tight-binding parameters
So far our calcuations have employed a minimal model with only $\gamma_{0}$, $\gamma_{1}$, and $U$, Eq. (1). To assess the influence of additional tight-binding parameters, we consider three examples: Rhombohedral graphite (ABC stacking), hexagonal boron nitride [27] (AA stacking), and hexagonal boron phosphide [60, 61, 62] (ABA stacking).
X.1 ABC stacking: rhombohedral graphite
We consider a tight-binding model of rhombohedral graphite containing skew interlayer hopping $\gamma_{3}$ and $\gamma_{4}$, and next-nearest layer hopping $\gamma_{2}$. In a basis of a single orbital on each atomic site $(A_{1},B_{1},A_{2},B_{2},\ldots,A_{N},B_{N})$, the Hamiltonian (1) is modified [63] as
$H=\begin{pmatrix}D_{1}&V&W&0&0&\ldots\\ V^{\dagger}&D_{2}&V&W&0&\ldots\\ W^{\dagger}&V^{\dagger}&D_{3}&V&W&\ldots\\ 0&W^{\dagger}&V^{\dagger}&D_{4}&V&\ldots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix},$ (43)
where intralayer blocks $D_{i}$ are defined in Eq. (2) and we use $U_{i}=U$ for all $i$ (for a faultless system). Interlayer blocks are
$V$ $=$ $\begin{pmatrix}\gamma_{4}f({\bf q})&-\gamma_{3}f^{\ast}({\bf q})\\ \gamma_{1}&\gamma_{4}f({\bf q})\end{pmatrix},$ (44)
and next-nearest layer blocks are
$W$ $=$ $\begin{pmatrix}0&\gamma_{2}/2\\ 0&0\end{pmatrix}.$ (45)
The band structures and corresponding conductivities for this model are shown in Fig. 17(e)-(h). We use the following parameter values [40, 63], $\gamma_{0}=3.16\,$eV, $\gamma_{1}=0.381\,$eV, $\gamma_{2}=-0.02\,$eV, $\gamma_{3}=0.315\,$eV, $\gamma_{4}=0.044\,$eV with $U=0.1\,$eV. Fig. 17(e) and (f) are for the faultless system, whereas Fig. 17(g) and (h) are for an atomically-sharp soliton at the center of the system. In comparison to the minimal model, it can be seen that the additional parameters have a generally small effect on the band structure and conductivity, without producing any qualitative changes.
Figure 18: (a), (b) band structure of hexagonal boron nitride with $N=10$ layers and AA stacking, plotted across the whole Brillouin zone with (a) a faultless system, and (b) a soliton at the center. Plots are obtained by diagonalizing Hamiltonian (46) with parameters $\gamma_{0}=2.33\,$eV, $\gamma_{1}=0.5\,$eV, $\gamma_{n}=-0.4\,$eV, $\gamma_{2}=-0.1\,$eV, and $U=3.625\,$eV [64]. (c), (d) band structure of hexagonal boron phosphide with $N=10$ layers and ABA stacking, plotted across the whole Brillouin zone with (c) a faultless system, and (d) a soliton at the center. Plots are obtained by diagonalizing Hamiltonian (49) with parameters $\gamma_{0}=1.65\,$eV, $\gamma_{1}=0.761\,$eV, $\gamma_{3}=-0.456\,$eV, $\gamma_{4}=0.255\,$eV, and $U=0.636\,$eV [61].
X.2 AA stacking: hexagonal boron nitride
As a second example of the influence of additional tight-binding parameters, we use a tight-binding model of AA-stacked hexagonal boron nitride (h-BN) with parameters fitted to calculations using density functional theory (DFT) and the GW approximation for a faultless system [64]. In this model [64], tight-binding parameters are $\gamma_{0}=2.33\,$eV, $\gamma_{1}=0.5\,$eV, and $U=3.625\,$eV. Additional parameters, as compared to the minimal model, are in-plane next-nearest neighbor hopping $\gamma_{n}=-0.4\,$eV and next-nearest layer hopping $\gamma_{2}=-0.1\,$eV.
In this tight-binding model, for AA-stacking, the Hamiltonian (1) is modified [64] as
$H=\begin{pmatrix}\tilde{D}_{1}&\tilde{V}&\tilde{W}&0&0&\ldots\\ \tilde{V}^{\dagger}&\tilde{D}_{2}&\tilde{V}&\tilde{W}&0&\ldots\\ \tilde{W}^{\dagger}&\tilde{V}^{\dagger}&\tilde{D}_{3}&\tilde{V}&\tilde{W}&\ldots\\ 0&\tilde{W}^{\dagger}&\tilde{V}^{\dagger}&\tilde{D}_{4}&\tilde{V}&\ldots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix},$ (46)
written in terms of $2\times 2$ blocks. Intralayer blocks are
$\tilde{D}_{i}=\begin{pmatrix}U_{i}-\gamma_{n}|f({\bf q})|^{2}&-\gamma_{0}f({\bf q})\\ -\gamma_{0}f^{\ast}\!({\bf q})&-U_{i}-\gamma_{n}|f({\bf q})|^{2}\end{pmatrix},$ (47)
where $U_{i}=U$ for odd $i$ and $U_{i}=-U$ for even $i$ (for a faultless system). Interlayer blocks are
$\tilde{V}=\begin{pmatrix}\gamma_{1}&0\\ 0&\gamma_{1}\end{pmatrix};\qquad\tilde{W}=\begin{pmatrix}\gamma_{2}&0\\ 0&\gamma_{2}\end{pmatrix}.$ (48)
The band structure plotted across the whole Brillouin zone is shown for a faultless system in Fig. 18(a) and for an atomically-sharp soliton at the center in Fig. 18(b). For the faultless system, the band gap is near the K point but is indirect in the presence of additional parameters, Fig. 18(a). For the soliton, there are two states localized on the soliton within the bulk band gap, Fig. 18(b), and this is qualitatively similar to the minimal model, Fig. 11(e). However, h-BN is in the wide band gap regime with $U/\gamma_{1}\approx 7.25$, and, as a result, the soliton states do not extend very far into the band gap, in relative terms.
X.3 ABA stacking: hexagonal boron phosphide
As a third example of the influence of additional tight-binding parameters, we use a tight-binding model of ABA-stacked hexagonal boron phosphide (h-BP) with parameters fitted to calculations using DFT for a faultless system [61]. As compared to the minimal model, we add skew interlayer couplings $\gamma_{3}$ and $\gamma_{4}$. In this tight-binding model, for ABA-stacking, the Hamiltonian (1) is modified [65] as
$H=\begin{pmatrix}D_{1}&V&0&0&0&\ldots\\ V^{\dagger}&D_{2}&V^{\dagger}&0&0&\ldots\\ 0&V&D_{3}&V&0&\ldots\\ 0&0&V^{\dagger}&D_{4}&V^{\dagger}&\ldots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix},$ (49)
where intralayer blocks $D_{i}$ are defined in Eq. (2) and we use $U_{i}=U$ for all $i$ (for a faultless system). Interlayer blocks $V$ are given by Eq. (44). We use parameter values fit to DFT [61] near the K point for bilayers, neglecting small differences in the hopping between boron and phosphorous atoms, giving $\gamma_{0}=1.65\,$eV, $\gamma_{1}=0.761\,$eV, $\gamma_{3}=-0.456\,$eV, $\gamma_{4}=0.255\,$eV, and $U=0.636\,$eV.
The band structure plotted across the whole Brillouin zone is shown for a faultless system in Fig. 18(c) and for an atomically-sharp soliton at the center in Fig. 18(d). For the faultless system, the band gap is near the K point but is indirect in the presence of additional parameters, Fig. 18(c). For the soliton, there is a state localized on the soliton within the bulk band gap, above the bulk valence bands, Fig. 18(d). This is qualitatively similar to the minimal model, Fig. 7(c). Note that h-BP is in a regime of a relatively-small band gap with $U/\gamma_{1}\approx 0.84$.
XI Relevance to specific materials
@@
Our results apply to binary compounds with a honeycomb structure. A well-studied example is h-BN [27] which occurs predominantly with AA stacking [71, 72, 74, 73, 75, 30, 31, 29, 76, 77], although samples with ABA [78, 31, 79, 80, 29, 81, 82] and ABC stacking [83, 84, 85, 86, 87] have also been fabricated. Density functional theory [28, 38, 30, 29] predicts that ABA stacking has a total energy comparable to AA, but that ABC stacking is less energetically favorable. Excitons localized on structural defects in h-BN have been discussed in the context of observed optical properties [51, 52, 53, 54], and features in X-ray spectroscopy have been attributed to interlayer stacking faults [55]. Domain walls in the in-plane direction have been observed experimentally [31, 56, 57] and discussed theoretically [88], and the nonlinear optical properties of multilayers with a controllable twist in the vertical (stacking) direction have been measured and analyzed [89].
The stacking types to be realized experimentally (AA, ABA and ABC) all have different atoms connected in the vertical direction, i.e., boron to nitrogen, and a fault consisting of a connection between two identical atoms, i.e., boron to boron or nitrogen to nitrogen, will be energetically expensive. Single layer defects, as described in Section VII, were modeled using DFT [38] for AA, AA, ABA and ABA stacking, and it was found that the formation energy of such faults is very small. For example, an AA-type defect in AA stacking was estimated to cost of the order of 505050\,50meV. Even so, it is likely that such a defect is difficult to realize experimentally because a relative shift in the in-plane direction would yield the more favorable ABA stacking [78, 29]. Nevertheless, experiments [90, 91] have recently demonstrated the possiblilty of engineering stacking of boron nitride (to create ferroelectric materials), and a relative twist between layers creates domains of ABA stacking separated, in the in-plane direction, by domain walls and regions of AA stacking.
@@
XII Conclusions
Using a minimal tight-binding model, we determined the electronic properties of thin films of binary compounds with two atoms per unit cell arranged as stacked two-dimensional honeycomb lattices. The two atoms per cell are assigned different onsite energies. We considered six different stacking orders to determine whether a fault in the texture of onsite energies in the stacking direction supports localized states. Faults for ABC, ABA, and AA stacking support localized states within the band gap, whereas faults for ABC, ABA, and AA stacking do not. For ABC and ABA stacking, there is a single localized state within the bulk band gap and, for ABC stacking, this state may hybridize with surface states. For AA stacking, there are two states within the bulk band gap which may hybridize with each other, depending on parameter values.
We consider smooth solitons where the texture of onsite energies changes over a length scale greater than the interlayer spacing, leading to different band structures as compared to atomically sharp solitons. In particular, a smooth soliton in AA stacking results in a flat band at zero energy and a corresponding narrow peak in the density of states. We show that this feature is fairly robust to long-range correlated disorder. Finally, we also consider the band structure due to single layer defects [38] where the signs of the onsite energies are reversed on a single layer only without a subsequent change of texture. Overall, different stackings and types of fault produce a range of contrasting electronic band structures and densities of states which will manifest themselves in various transport and spectroscopic measurements [51, 52, 53, 54, 55, 31, 56, 57].
@@ Acknowledgements.
Appendix A Transport calculation
Here we provide details of the conductivity calculation for rhombohedral stacking shown in Fig. 17. We consider a two-contact system with a central sample at $0<x<L$ and two semi-infinite leads on the left ($x<0$) and on the right ($x>L$). Following [58, 42], the leads are modeled by considering the system with a very large onsite energy, creating a large density of states. Wave matching is performed to match states in the central sample to those in the leads at the same $E$ and $k_{y}$ values.
We consider the continuum limit near the valley at wave vector ${\mathbf{K_{+}}}=(4\pi/(3a),0)$ so that $f({\mathbf{K_{+}}}+{\bf k})\approx-\sqrt{3}a(k_{x}-ik_{y})/2$. Following [63], we write the Hamiltonian (43) as
$H=\begin{pmatrix}D_{1}&V&W&0&0&\ldots\\ V^{\dagger}&D_{2}&V&W&0&\ldots\\ W^{\dagger}&V^{\dagger}&D_{3}&V&W&\ldots\\ 0&W^{\dagger}&V^{\dagger}&D_{4}&V&\ldots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix},$ (50)
where
$D_{j}=\begin{pmatrix}U_{j,A}&v\pi^{\dagger}\\ v\pi&U_{j,B}\end{pmatrix};\quad V=\begin{pmatrix}-v_{4}\pi^{\dagger}&v_{3}\pi\\ \gamma_{1}&-v_{4}\pi^{\dagger}\end{pmatrix},$ (51)
with $\pi=\hbar(k_{x}+ik_{y})$, $\pi^{\dagger}=\hbar(k_{x}-ik_{y})$, $v_{3}=\sqrt{3}a\gamma_{3}/(2\hbar)$, and $v_{4}=\sqrt{3}a\gamma_{4}/(2\hbar)$. Next-nearest layer hopping is described by
$W=\begin{pmatrix}0&\gamma_{2}/2\\ 0&0\end{pmatrix}.$ (52)
In the center of the system, $0<x<L$, we set $U_{A,j}=-U_{B,j}=U_{j}$ for all $j$ (for a faultless system) and numerically determine $2N$ eigenstates $\Psi_{j}$ of the Hamiltonian (50) each with a value of $k_{x}$ denoted $k_{j}$ where $j=1,2,\ldots,2N$. Then the states in the center can be written as
$\psi^{\ell}=\sum_{j}c_{j}^{\ell}\Psi_{j}e^{ik_{j}x+ik_{y}y},$ (53)
where index $\ell=1,2,\ldots,N$ denotes $N$ different incoming states.
In the contacts, we set $U_{A,j}=U_{B,j}=-U_{\infty}$ for all $j$, where $U_{\infty}\gg\gamma_{1},E$. Then we numerically determine the eigenstates of the Hamiltonian (50). There are $N$ right-moving solutions $\Upsilon_{\ell}$ with $k_{x}=Q_{\ell}^{(R)}$ where $\mathrm{Re}(Q_{\ell}^{(R)})>0$, and three left-moving solutions $\Phi_{\ell}$ with $k_{x}=-Q_{\ell}^{(L)}$ where $\mathrm{Re}(Q_{\ell}^{(L)})>0$, where $\ell=1,2,\ldots,N$. Each of them is normalized to carry unit flux by setting $\Upsilon_{\ell}^{\dagger}S_{x}\Upsilon_{\ell}=\Phi_{\ell}^{\dagger}S_{x}\Phi_{\ell}=1$ where $S_{x}$ is proportional to the current operator,
$S_{x}=\begin{pmatrix}0&1&0&0&0&0&\ldots\\ 1&0&0&0&0&0&\ldots\\ 0&0&0&1&0&0&\ldots\\ 0&0&1&0&0&0&\ldots\\ 0&0&0&0&0&1&\ldots\\ 0&0&0&0&1&0&\ldots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix},$ (54)
In the left contact at $x=0$, there are $N$ possible states, $\ell=1,2,\ldots,N$, describing incoming flux,
${\cal L}^{\ell}=\Upsilon_{\ell}e^{ik_{y}y}+\sum_{j}r_{j}^{\ell}\Phi_{j}e^{ik_{y}y},$ (55)
where $r_{j}^{\ell}$ is a reflection amplitude from state ${\ell}$ to state $j$. In the right contact at $x=L$, there are $N$ corresponding transmitted states:
${\cal R}^{\ell}=\sum_{j}t_{j}^{\ell}\Upsilon_{j}e^{iQ_{j}^{(R)}L+ik_{y}y},$ (56)
where $t_{j}^{\ell}$ is a transmission amplitude from state ${\ell}$ to state $j$.
We consider wave-matching $\psi^{\ell}={\cal L}^{\ell}$ at $x=0$ and $\psi^{\ell}={\cal R}^{\ell}$ at $x=L$ for $\ell=1,2,\ldots,N$. For each $\ell$ value, there are $4N$ unknowns consisting of $2N$ values of $c_{j}^{\ell}$, $N$ values of $r_{j}^{\ell}$, and $N$ values of $t_{j}^{\ell}$, and there are $4N$ boundary conditions due to continuity of the wave function components. Following [42], the matching conditions for all $\ell$ values may be expressed as $MC=A$ where $C$ is a $4N\times N$ matrix of unknown coefficients
$C=\begin{pmatrix}\hat{r}^{T}\\ \hat{c}\\ \hat{t}^{T}\end{pmatrix},$ (57)
where $\hat{r}_{j\ell}=r_{\ell}^{j}$ is $N\times N$, $\hat{c}_{j\ell}=c_{j}^{\ell}$ is $2N\times N$, and $\hat{t}_{j\ell}=t_{\ell}^{j}$ is $N\times N$. $A$ is a $4N\times N$ matrix of incoming states,
$A=\begin{pmatrix}-\Upsilon_{1}&-\Upsilon_{2}&-\Upsilon_{3}&\ldots&-\Upsilon_{N}\\ \hat{0}&\hat{0}&\hat{0}&\ldots&\hat{0}\end{pmatrix},$ (58)
where $\hat{0}$ represents a $2N\times 1$ column vector of zeros. $M$ is a $4N\times 4N$ matrix,
[…]
where $z_{j}=e^{ik_{j}L}$ for $j=1,2,\ldots,2N$, and $y_{\ell}=e^{iQ_{\ell}^{(R)}L}$ for $\ell=1,2,\ldots,N$. In the leads with $U_{\infty}\gg\gamma_{1},E$, the dispersion is approximately linear so that $U_{\infty}\approx\hbar vQ_{\ell}^{(R)}$. For our numerical calculations, Fig. 17, we choose $U_{\infty}=50\gamma_{1}$ so that $Q_{\ell}^{(R)}\approx 50k_{c}$.
The unknown coefficients are found by determining $C=M^{-1}A$. Then, for a given energy value, the conductivity [59, 42] is given by
$\sigma(E)=\frac{4e^{2}}{h}\frac{L}{W}\mathrm{tr}(\hat{t}\hat{t}^{\dagger}),$ (60)
where $L$ is the length and $W$ is the width of the sample. The trace may evaluated by summing over the eigenvalues of $\hat{t}\hat{t}^{\dagger}$ for all $k_{y}$ values where, for periodic boundary conditions [42], $k_{y}=2\pi n/W$ for integer $n=0,\pm 1,\pm 2,\ldots$.